# Regression of Suspension Violin Modes in KAGRA O3GK Data with Kalman Filters

### v2

@@ IV.3 Thermal Noise Subtraction using O3GK KAGRA Data
Table 2: Kalman bands used to suppress suspension thermal peaks including central frequency $f_{c}$ and bandwidth $\Delta f_{c}$.
@@ IV.4 Test Injection into O3GK KAGRA Data
In order to verify that gravitational-wave signals are preserved, simulated signals must be injected into the strain channel prior to filtering, and the recovered waveforms should be compared after subtraction. From the perspective of gravitational-wave detection, the precise evaluation of any residual contamination is critical and requires further quantification.
@@
For the computation of the signal-to-noise ratio (SNR), denoted $\rho$, as defined in Eq.(22), a reference power spectral density (PSD) is estimated from a 4096,s segment of O3GK data.
@@ Appendix A Resonant Frequency Extractions
The values used in Kalman filtering are shown in Table. 3. The resonant frequencies for the 3 violin modes are taken from the PTEP [2], the other parameters ($Q_{i}$ values and $G_{i}$ values) used in the Lorentzian profile results from a fit. The ${\rm FWHM}_{i}$ are also shown.
@@ Appendix B Derivation of the second-order resonance equations
Consider a mirror of mass $M$ suspended by four fibers at positions $(x_{i},y_{i},z_{i})$ relative to the mirror’s center of mass (CoM). The mirror has moments of inertia $I_{y}$ and $I_{z}$ about the $y$ (pitch) and $z$ (yaw) axes, respectively. Roll effects are neglected, as they have negligible influence on the measurable interferometer output signal. Each fiber has a mass $m_{i}$ and stiffness $k_{i}$. The beam propagates along $x$, with $y$ horizontal and $z$ vertical. The relative elongation of the $i$-th fiber is defined as
$X_{i}(t)=x_{i}(t)-X_{0,i}(t),$ (23)
@@
$\left\{\begin{aligned} M\ddot{X}&+\sum_{i=1}^{4}F_{i,x}(t)=0,\\ I_{z}\ddot{\theta}_{z}&-\sum_{i=1}^{4}y_{i}\,F_{i,x}(t)=0,\\ I_{y}\ddot{\theta}_{y}&+\sum_{i=1}^{4}z_{i}\,F_{i,x}(t)=0.\end{aligned}\right$. (25)
@@
$\tilde{X}_{i}(\omega)=\frac{m_{i}\omega^{2}}{k_{i}-m_{i}\omega^{2}}\,\Big[\tilde{X}(\omega)-y_{i}\tilde{\theta}_{z}(\omega)+z_{i}\tilde{\theta}_{y}(\omega)\Big].$ (27)
@@
The mirror equations of motion in Fourier domain take the form
$\left\{\begin{aligned} -\,M\omega^{2}\tilde{X}&+\sum\tilde{K}_{i}(\omega)\,~~~\bigl[\tilde{X}-y_{i}\tilde{\theta}_{z}+z_{i}\tilde{\theta}_{y}\bigr]&=0,\\[4.0pt] -\,I_{z}\omega^{2}\tilde{\theta}_{z}&-\sum y_{i}\,\tilde{K}_{i}(\omega)\,\bigl[\tilde{X}-y_{i}\tilde{\theta}_{z}+z_{i}\tilde{\theta}_{y}\bigr]&=0,\\[4.0pt] -\,I_{y}\omega^{2}\tilde{\theta}_{y}&+\sum z_{i}\,\tilde{K}_{i}(\omega)\,\bigl[\tilde{X}-y_{i}\tilde{\theta}_{z}+z_{i}\tilde{\theta}_{y}\bigr]&=0.\end{aligned}\right$.
@@
$\underbrace{\begin{bmatrix}M&0&0\\ 0&I_{z}&0\\ 0&0&I_{y}\end{bmatrix}}_{\mathbf{M}}\begin{bmatrix}-\,\omega^{2}\tilde{X}\\ -\,\omega^{2}\tilde{\theta}_{z}\\ -\,\omega^{2}\tilde{\theta}_{y}\end{bmatrix}\;+\;\underbrace{\begin{bmatrix}K_{00}&0&0\\ 0&K_{yy}&0\\ 0&0&K_{zz}\end{bmatrix}}_{\mathbf{K}(\omega)}\begin{bmatrix}\tilde{X}\\[4.0pt] \tilde{\theta}_{z}\\[4.0pt] \tilde{\theta}_{y}\end{bmatrix}=\mathbf{0},$ (30)
@@
$K_{yy}(\omega)$ $=\sum_{i}\tilde{K}_{i}(\omega)y_{i}^{2},\qquad K_{zz}(\omega)=\sum_{i}\tilde{K}_{i}(\omega)z_{i}^{2}.$
@@
$\det\!\left(\mathbf{K}(\omega)-\omega^{2}\mathbf{M}\right)=0.$
@@
By contrast, antisymmetric oscillations—such as pairs of fibers stretching in opposite directions—couple directly to the mirror’s rotational degrees of freedom. For simplicity purpose, we are also assuming identical fibers ($k_{i}=k$, $m_{i}=m$). In this case, the angular eigenvalues describe the yaw and pitch sidebands as
$\Delta f_{\rm yaw}\;\approx\;\frac{f_{0}}{2I_{z}}\sum_{i=1}^{N}m\,y_{i}^{2},\qquad\Delta f_{\rm pitch}\;\approx\;\frac{f_{0}}{2I_{y}}\sum_{i=1}^{N}m\,z_{i}^{2}.$
@@
$I_{x}=\dfrac{1}{2}MR^{2},\qquad I_{y}=I_{z}=\dfrac{M}{12}\left(3R^{2}+L^{2}\right),$ (31)
Typically, $d_{y}$ is of order the mirror radius, while $d_{z}\simeq 0$ when the fibers lie in the horizontal plane. For sapphire fibers and cylindrical test masses, this scaling gives $\Delta f_{\rm yaw}\approx 0.10~\text{Hz}$. Since the actual suspension is not perfectly symmetric (bonding, stiffness, clamp anisotropy, small tension imbalances), this value should be viewed as a lower bound; realistic asymmetries can increase $\Delta f_{\rm yaw}$ into the few-tenths-of-hertz range. Each suspension fiber supports nearly two degenerate violin polarizations: a x-direction mode that couples strongly to the interferometer’s arm length readout, and a y-axis companion that couples weakly through asymmetries and miscentering. Both modes share the same intrinsic Q, but the yaw-like resonance appears at lower amplitude and slightly shifted frequency, producing the observed doublets in the violin-mode spectrum.
@@ Appendix C Process & Measurement Noise Formula
$\ddot{x}(t)+\frac{\omega_{0}}{Q}\,\dot{x}(t)+\omega_{0}^{2}x(t)=\frac{1}{m}\,\xi_{x}(t),$ (32)
@@
$\left\{\begin{aligned} \langle\xi_{x}(t)\,\xi_{x}(t^{\prime})\rangle&=S_{F}^{(2)}(\omega_{0})\,\delta(t-t^{\prime}),\\[1.72218pt] S_{F}^{(2)}(\omega_{0})&=2k_{B}T\,\mathrm{Re}\left[\,Z(\omega_{0})\right]\;=\;2k_{B}T\,m\,\frac{\omega_{0}}{Q}\end{aligned}\right$. (33)
@@
$\langle\xi_{x}(t)\,\xi_{x}(t^{\prime})\rangle=0\quad\text{for }t\neq t^{\prime},$
@@
$=\frac{k_{B}T}{2}\,\frac{\omega_{0}}{Q\,m\,\omega_{c}^{2}}\,\delta(t-t^{\prime})\,\mathbf{I}_{2}.$ (34)
@@
$\sigma_{w,i}^{2}=\frac{k_{B}T}{4\Delta f_{c}}\,\frac{\omega_{0}}{Q\,m\,\omega_{c}^{2}}.$ (35)
@@
$\sigma_{v}^{2}=\frac{1}{\Delta f_{c}}\,\frac{\bar{P}_{m}}{2}.$ (36)
Figure 15: (a) Top: LIGO 40m PSD. The colored vertical lines shows the violin modes frequencies for each of the 8th first modes. (b), (c) and (d) Respectively, the 1st, 2nd and 4th violin modes in LIGO 40m.

### v3

@@ IV.3 Thermal Noise Subtraction using O3GK KAGRA Data
Table 2: Kalman bands used to suppress suspension thermal peaks, including central frequency $f_{c}$ and bandwidth $\Delta f_{c}$.
@@ IV.4 Test Injection into O3GK KAGRA Data
To verify that gravitational-wave signals are preserved, simulated signals must be injected into the strain channel prior to filtering, and the recovered waveforms should be compared after subtraction. From the perspective of gravitational-wave detection, the precise evaluation of any residual contamination is critical and requires further quantification.
@@
For the computation of the signal-to-noise ratio (SNR), denoted $\rho$, as defined in Eq.(22), a reference power spectral density (PSD) is estimated from a 4096 s segment of O3GK data.
@@ Appendix A Resonant Frequency Extractions
The values used in Kalman filtering are shown in Table. 3. The resonant frequencies for the three violin modes are taken from the PTEP [2], other parameters ($Q_{i}$ values and $G_{i}$ values) used in the Lorentzian profile result from a fit. The ${\rm FWHM}_{i}$ are also shown.
@@ Appendix B Derivation of the second-order resonance equations
Consider a mirror of mass $M$ suspended by four fibers at positions $(x_{i},y_{i},z_{i})$ relative to the mirror’s center of mass (CoM). The mirror has moments of inertia $I_{y}$ and $I_{z}$ about the $y$ (pitch) and $z$ (yaw) axes, respectively. Roll effects are neglected, as they have a negligible influence on the measurable interferometer output signal. Each fiber has a mass $m_{i}$ and stiffness $k_{i}$. The beam propagates along the $x$ direction, with $y$ being horizontal and $z$ being vertical. The relative elongation of the $i$-th fiber is defined as
$X_{i}(t)=x_{i}(t)-X_{0,i}(t)$ (23)
@@
$\left\{\begin{aligned} M\ddot{X}&+\sum_{i=1}^{4}F_{i,x}(t)=0\\ I_{z}\ddot{\theta}_{z}&-\sum_{i=1}^{4}y_{i}\,F_{i,x}(t)=0\\ I_{y}\ddot{\theta}_{y}&+\sum_{i=1}^{4}z_{i}\,F_{i,x}(t)=0\end{aligned}\right$. (25)
@@
$\tilde{X}_{i}(\omega)=\frac{m_{i}\omega^{2}}{k_{i}-m_{i}\omega^{2}}\,\Big[\tilde{X}(\omega)-y_{i}\tilde{\theta}_{z}(\omega)+z_{i}\tilde{\theta}_{y}(\omega)\Big]$ (27)
@@
The mirror equations of motion in the Fourier domain take the form:
$\left\{\begin{aligned} -\,M\omega^{2}\tilde{X}&+\sum\tilde{K}_{i}(\omega)\,~~~\bigl[\tilde{X}-y_{i}\tilde{\theta}_{z}+z_{i}\tilde{\theta}_{y}\bigr]&=0\\[4.0pt] -\,I_{z}\omega^{2}\tilde{\theta}_{z}&-\sum y_{i}\,\tilde{K}_{i}(\omega)\,\bigl[\tilde{X}-y_{i}\tilde{\theta}_{z}+z_{i}\tilde{\theta}_{y}\bigr]&=0\\[4.0pt] -\,I_{y}\omega^{2}\tilde{\theta}_{y}&+\sum z_{i}\,\tilde{K}_{i}(\omega)\,\bigl[\tilde{X}-y_{i}\tilde{\theta}_{z}+z_{i}\tilde{\theta}_{y}\bigr]&=0\end{aligned}\right$.
@@
$\underbrace{\begin{bmatrix}M&0&0\\ 0&I_{z}&0\\ 0&0&I_{y}\end{bmatrix}}_{\mathbf{M}}\begin{bmatrix}-\,\omega^{2}\tilde{X}\\ -\,\omega^{2}\tilde{\theta}_{z}\\ -\,\omega^{2}\tilde{\theta}_{y}\end{bmatrix}\;+\;\underbrace{\begin{bmatrix}K_{00}&0&0\\ 0&K_{yy}&0\\ 0&0&K_{zz}\end{bmatrix}}_{\mathbf{K}(\omega)}\begin{bmatrix}\tilde{X}\\[4.0pt] \tilde{\theta}_{z}\\[4.0pt] \tilde{\theta}_{y}\end{bmatrix}=\mathbf{0}$ (30)
@@
$K_{yy}(\omega)$ $=\sum_{i}\tilde{K}_{i}(\omega)y_{i}^{2},\qquad K_{zz}(\omega)=\sum_{i}\tilde{K}_{i}(\omega)z_{i}^{2}$
@@
$\det\!\left(\mathbf{K}(\omega)-\omega^{2}\mathbf{M}\right)=0$
@@
By contrast, antisymmetric oscillations—such as pairs of fibers stretching in opposite directions—couple directly to the mirror’s rotational degrees of freedom. For simplicity purposes, we are also assuming identical fibers ($k_{i}=k$, $m_{i}=m$). In this case, the angular eigenvalues describe the yaw and pitch sidebands as
$\Delta f_{\rm yaw}\;\approx\;\frac{f_{0}}{2I_{z}}\sum_{i=1}^{N}m\,y_{i}^{2},\qquad\Delta f_{\rm pitch}\;\approx\;\frac{f_{0}}{2I_{y}}\sum_{i=1}^{N}m\,z_{i}^{2}$
@@
$I_{x}=\dfrac{1}{2}MR^{2},\qquad I_{y}=I_{z}=\dfrac{M}{12}\left(3R^{2}+L^{2}\right)$ (31)
Typically, $d_{y}$ is of order the mirror radius, while $d_{z}\simeq 0$ when the fibers lie in the horizontal plane. For sapphire fibers and cylindrical test masses, this scaling gives $\Delta f_{\rm yaw}\approx 0.10~\text{Hz}$. Since the actual suspension is not perfectly symmetric (bonding, stiffness, clamp anisotropy, small tension imbalances), this value should be viewed as a lower bound; realistic asymmetries can increase $\Delta f_{\rm yaw}$ into the few-tenths-of-hertz range. Each suspension fiber supports nearly two degenerate violin polarizations: an x-direction mode that couples strongly to the interferometer’s arm length readout, and a y-axis companion that couples weakly through asymmetries and miscentering. Both modes share the same intrinsic Q, but the yaw-like resonance appears at a lower amplitude and slightly shifted frequency, producing the observed doublets in the violin-mode spectrum.
@@ Appendix C Process & Measurement Noise Formula
$\ddot{x}(t)+\frac{\omega_{0}}{Q}\,\dot{x}(t)+\omega_{0}^{2}x(t)=\frac{1}{m}\,\xi_{x}(t)$ (32)
@@
$\left\{\begin{aligned} \langle\xi_{x}(t)\,\xi_{x}(t^{\prime})\rangle&=S_{F}^{(2)}(\omega_{0})\,\delta(t-t^{\prime})\\[1.72218pt] S_{F}^{(2)}(\omega_{0})&=2k_{B}T\,\mathrm{Re}\left[\,Z(\omega_{0})\right]\;=\;2k_{B}T\,m\,\frac{\omega_{0}}{Q}\end{aligned}\right$. (33)
@@
[…]
@@
$=\frac{k_{B}T}{2}\,\frac{\omega_{0}}{Q\,m\,\omega_{c}^{2}}\,\delta(t-t^{\prime})\,\mathbf{I}_{2}$ (34)
@@
$\sigma_{w,i}^{2}=\frac{k_{B}T}{4\Delta f_{c}}\,\frac{\omega_{0}}{Q\,m\,\omega_{c}^{2}}$ (35)
@@
$\sigma_{v}^{2}=\frac{1}{\Delta f_{c}}\,\frac{\bar{P}_{m}}{2}$ (36)
Figure 15: (a) Top: LIGO 40m PSD. The colored vertical lines show the violin modes frequencies for each of the 8th first modes. (b), (c) and (d) Respectively, the 1st, 2nd and 4th violin modes in LIGO 40m.